p-2/3 +p/4 =1/2
4p-8/12 + 3p/12 =6
7p-2/3=6
lost, lost lost

Answer by sdmmadam@yahoo.com(530)

1)p-2/3 +p/4 =1/2 ----(1)
p+p/4 =1/2+ 2/3 (grouping like terms changing sign while changing side)
Multiplying through out by 12 (which is the lcm of 2,3 and 4
12p +3p = 6 + 8
15p = 14
p=14/15
Answer: p = 14/15
Verification: Putting p=14/15 in (1),LHS simplifies to 1/2 = RHS
2)4p-8/12 + 3p/12 =6 ----(2)
Multiplying through out by 12
4pX12-8+3p =6X12
48p-8+3p=72
(48p+3p) = 72+8
51p = 80
p= 80/51
Answer:p= 80/51
Verification: Putting p=80/51 in (2),
LHS simplifies to 320/51-2/3+20/51 = 306/51 =6 = RHS
3)7p-2/3=6 ----(3)
7p = 6+2/3
7p =(6X3+2)/3
7p = 20/3
Dividing by 7 (that is multiplying by (1/7)
p= (20/3)X(1/7) = 20/21
Answer: p = 20/21
Verification: p = 20/21 in LHS of (1)
gives 7p-2/3 = 7X(20/21)-2/3 = 20/3-2/3 = (20-2)/3 = 18/3 = 6 =RHS